OpenStudy (anonymous):

Lecture #3 textbook homework check. The non-colored problems do not have answers at the back of the book, so I would like to swap answers chapter 6, #46 and #54. (Other people may post answers to the non-colored chapter 1 problems, but I did those a while back and I did't want to redo them.)

OpenStudy (anonymous):

A&E Chapter 6, #46) \[\nu=c/\lambda=c/3.21m=93.4MHz\]\[\lambda=c/\nu=c/82MHz=3.66m\]\[\lambda=c/\nu=c/112MHz=2.68m\] A&E Chapter 6, #54) \[E=h\nu\]\[Photons=Output/E=Output/h\nu\]\[=50kW/(h*100.7MHz)=7.493*10^{32}photon/s\]

OpenStudy (anonymous):

Please note that my answers are not necessarily correct! (I want to swap answers to see if I made a mistake or not.)